The Royal Fruit Company produces two types of fruit drinks. The first type is
60%
pure fruit juice, and the second type is
85%
pure fruit juice. The company is attempting to produce a fruit drink that contains
80%
pure fruit juice. How many pints of each of the two existing types of drink must be used to make
180
pints of a mixture that is
80%
pure fruit juice?

Answer by TimothyLamb(4379)

x = pints of 60%
y = pints of 85%
---
x + y = 180
60x + 85y = 80*180
---
put the system of linear equations into standard form
---
x + y = 180
60x + 85y = 14400
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = pints of 60% = 36
y = pints of 85% = 144
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---